OpenStudy (anonymous):

On the graph of the equation 2x + 5y = -30, what is the value of the y-intercept? 15 -15 6 -6

OpenStudy (abmon98):

5y=-2x-30 y=-2/5x-6

OpenStudy (abmon98):

Y=mx+c C is the y-intercept

OpenStudy (the_fizicx99):

Solve for "y", 2x + 5y = -30 5y = -2x - 30 y = -2/5x - 6 the y intercept is the constant. Which in this case it's -6